OpenStudy (anonymous):

HELP ONLY 2 QUESTIONS! Simplify completely 25 w to the sixth power over 10 w to the third power divided by 30 w squared over 5 w 1 over 15 w to the fourth power 5 w squared over 12 12 over 5 w squared 15w4 Simplify completely quantity x squared minus 4 x plus 4 over quantity x squared plus 10 x plus 25 times quantity x plus 5 over quantity x squared plus 3 x minus 10 quantity x minus 2 over quantity x plus 5 x minus 2 all over the quantity x plus 5 end quantity squared the quantity x minus 2end quantity squared all over x plus 5

OpenStudy (whpalmer4):

\[\frac{\frac{25w^6}{10w^3}}{\frac{30w^2}{5w}}\]Is that the first expression? I would first do the fraction division, by multiplying the top half by the reciprocal of the bottom half: \[\frac{25w^6}{10w^3}*\frac{5w}{30w^2}\]Then use the rules of exponents to simplify, in particular \[\frac{x^n}{x^m} = x^{n-m}\]

OpenStudy (whpalmer4):

Try writing out the second problem with the Equation editor button. Make it easy for people to help you.